Sketch the level curves f−1(c) for the following functions:

f(x, y, z) = x − y2 − z2, c = −1, 0, 1.


1
Expert's answer
2020-10-20T18:29:03-0400

f(x, y, z) = x − y2 − z2, c = −1, 0.

1)c=-1

x − y2 − z2=-1

x+1=y2 +z2- elliptical paraboloid, formed by rotation around the x axis

2)c=0

x − y2 − z= 0

x= y2 +z2- elliptical paraboloid, formed by rotation around the x = axis

A graph is the family of elliptical paraboloids, where the top is shifts along the x-axis, depends on c-value.

3)c=1

x-1=y2 +z2- elliptical paraboloid, formed by rotation around the x axis